OpenStudy (anonymous):

A 210 ft. long rope is cut in three pieces. the first piece of the rope is 3 times as long as the second piece of rope The third piece of the rope is 2 times as long as the second piece of rope. What is the length of the longest piece of rope

OpenStudy (anonymous):

If you notice in this problem it says THE FIRST PIECE of the rope IS It says THE THIRD PIECE of the rope IS It never says THE SECOND PIECE of the rope IS.... SO the Second Piece gets X

OpenStudy (anonymous):

First piece = 3x Second piece = x Third piece = 2x Do you get this.. You read each sentence word for word

OpenStudy (anonymous):

Then you add the three pieces to get 210 3x + x + 2x = 210 6x = 210 x = 35

OpenStudy (anonymous):

Then substitute back in. First piece = 3x = 3(35) = 105 Second piece = x = 35 Third piece = 2x = 2(35) = 70

OpenStudy (anonymous):

Point A(-4,1) is in the standard (x,y) coordinate plane. what must be the coordinates of point B so that the line x=2 is the perpendicular bisector of line ab

OpenStudy (anonymous):

Point B can be anything with 1 as the y coordinate. x = 2 is a vertical line so you need something horizontal to (-4,1) which is anything with y = 1

OpenStudy (anonymous):

what is the slope of a line with the equation 2x+3y+6=0